OpenStudy (anonymous):

If sin theta = 1/2 and theta is in Quadrant 1, fins sin 2theta

OpenStudy (anonymous):

sin theta= 1/2 sin theta = sin pi/6 theta= pi/6 sin2theta= sin 2pi/6= sinpi/3= (square root of 3)/2

OpenStudy (cwrw238):

sin 2 theta = 2 sin theta cos theta cos^2 theta = 1 - sin^2 theta = 1 - (1/2)^2 = 3/4 cos theta = sqrt(3/4) = sqrt3 / 2 so sin 2 theta = 2 * 1/2 * sqrt3 / 2 = sqrt3 / 2

OpenStudy (anonymous):

thx

OpenStudy (cwrw238):

yw